OpenStudy (jiteshmeghwal9):

On solving \(\dfrac{25}{(x+y)}-\dfrac{3}{(x-y)}=1\), \(\dfrac{40}{(x+y)}+\dfrac{2}{(x-y)}=5\). we get :-

OpenStudy (anonymous):

Let a=1/(x+y) and b=1/(x-y)

OpenStudy (jiteshmeghwal9):

ok

OpenStudy (anonymous):

So, 25a-3b=1 40a-2b=5 Now, solve for a and b

Parth (parthkohli):

Nice solution!

OpenStudy (jiteshmeghwal9):

Thanx bro :)

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

And x=(1/a +1/b)/2